# Non-Holomorphic $A_{4}$ Modular Symmetry in Type-I Seesaw: Implications for Neutrino Masses and Leptogenesis

### v1

@@ 4.3 Implications to Neutrinoless Double Beta Decay:
Figure 7: Projected experimental sensitivity on effective Majorana mass parameter $m_{ee}$ for various neutrinoless double beta decay experiments. This image is initially discussed in ref [126] and utilized here to demonstrate the sensitivity of various neutrinoless double beta decay on the effective Majorana mass parameter $m_{ee}$. It can be seen that the Origin-X is the most promising beyond ton-scale experiment to probe the parameter space of $m_{ee}$ beyond 10 meV for normal ordering.
@@ 5 Leptogenesis within type-I seesaw framework
Leptogenesis has been an elegant framework for generating the observed baryon asymmetry of the Universe, which cannot be addressed within SM. The key idea of leptogenesis is to first create an asymmetry in the lepton sector, and subsequently the created lepton asymmetry becomes a net baryon asymmetry through anomalous $B+L$ violating sphaleron processes [31, 32, 33, 34]. In the original proposal of leptogenesis [23], the CP-violating and lepton number violating out of equilibrium decays of heavy Majorana right-handed neutrinos introduced within the Type-I seesaw mechanism. Many variants of leptogenesis [25, 26, 27, 131, 28] have been explored, and thermal leptogenesis is one of its kind, where the typical mass of right-handed neutrinos is considered at a high scale of around $10^{10}$  GeV GeV or more. Depending on how we account for flavor parts in the determination of CP-asymmetry, there are two classes of thermal leptogenesis: unflavored and flavored type, and we limit ourselves to unflavored leptogenesis in the rest of the analysis.
@@ 6 Conclusion
We have investigated a non-supersymmetric version of the Type I seesaw model in which the flavour structure is dictated by non-holomorphic modular forms of level three (polyharmonic Maass forms transforming under $\Gamma_{3}\simeq A_{4}$). The role of modular symmetry is to reduce the model parameters in the Yukawa sector by replacing conventional flavor fields with modular Yukawas. The structure of Yukawa couplings depends only on the complex modulus $\tau$; thereby, the mass matrices are controlled by $\tau$ and a few other minimal parameters. The framework simultaneously explains the observed pattern of lepton mixing and provides calculable predictions for observables that connect low-energy neutrino physics to high-scale leptogenesis.
@@
An important phenomenological implication of the present framework concerns neutrinoless double beta decay ($0\nu\beta\beta$), which serves as a key probe of the Majorana nature of neutrinos and lepton number violation. In our analysis, the estimated effective Majorana mass parameter $m_{ee}$ lies in the sub-meV to few-meV range, $m_{ee}\in[1.9\times 10^{-4},\,5.2\times 10^{-3}]$ eV, for normal ordering. This model predictions on effective Majorana mass parameter and half-lifes are well below the present experimental sensitivities from GERDA [144], KamLAND-Zen [145], and EXO-200 [146] which constrain $m_{ee}\lesssim(0.1-0.4)$ eV as well as the forthcoming ton-scale experiments such as CUPID [147], LEGEND-1000 [121], and nEXO [122], whose projected sensitivities extend down to $\mathcal{O}(0.006-0.027)$ eV. This covers the entire parameter space of $m_{1}-m_{ee}$ for Inverted Ordering and a very thin space in the upper region of the Normal Ordering . However, the limit of $m_{ee}$ that we obtained from our model can be accessed in the future beyond the ton-scale experiments like ORIGIN-X [124], THEIA[125], LEGEND-6000[123] etc as shown in Fig 7. Moreover, the tight parameter correlations imposed by the modular symmetry ensure that any future measurement of $0\nu\beta\beta$ would place stringent tests on the viability of this model, especially for the normal ordering scenario where the allowed region is already highly constrained.
@@ Acknowledgments
Appendix A Modular $A_{4}$ invariant Lagrangian
@@
Appendix B Yukawa couplings with different modular weights
@@
$\tau=x+iy,\quad x\equiv\operatorname{Re}[\tau],\quad y\equiv\operatorname{Im}[\tau],\quad q=e^{2\pi i\tau}.$ (B.1)
@@
$\hskip 56.9055pt+\frac{9q^{1/3}}{2\pi}\left(1+\frac{7q}{4}+\frac{8q^{2}}{7}+\frac{9q^{3}}{5}+\frac{14q^{4}}{13}+\frac{31q^{5}}{16}+\frac{20q^{6}}{19}+\dots\right),$ (B.2)
@@
$\hskip 56.9055pt-\frac{81q^{1/3}}{16\pi^{3}}\left(1+\frac{73q}{64}+\frac{344q^{2}}{343}+\frac{567q^{3}}{500}+\frac{20198q^{4}}{2197}+\frac{4681q^{5}}{4096}+\dots\right),$ (B.3)
@@
$M_{\ell}=v\,Y^{(0)}_{\ell}(\tau;\,\alpha_{\ell},\beta_{\ell},\gamma_{\ell}),\qquad M_{D}=v\,Y^{(-2)}_{D}(\tau;\,\alpha_{D},\beta_{D},\gamma_{D}),$ (B.4)
@@
Figure 11: We have schematically shown the Fundamental Domain for any data points of $Re[\tau]\;versus\;Im[\tau]$ satisfying the constraints of a Fundamental Domain in the upper half plane $\mathcal{H}$ .
@@
Table 10: Modular Yukawa triplets with three components evaluated at various representative complex moduli parameter $(\tau)$. These Yukawa couplings can be used to construct the charged-lepton and neutrino mass matrices by multiplying each triplet component by the corresponding complex prefactor (e.g. $(\alpha_{\ell},\beta_{\ell},\gamma_{\ell},\alpha_{D},\beta_{D},\gamma_{D})$) as described in the main text. The detailed discussion expansions listed in Appendix B.

### v2

@@ 4.3 Implications to Neutrinoless Double Beta Decay:
Figure 9: Projected experimental sensitivity on effective Majorana mass parameter $m_{ee}$ for various neutrinoless double beta decay experiments. This image is initially produced and discussed in ref [24] and utilized here to demonstrate the sensitivity of various neutrinoless double beta decay on the effective Majorana mass parameter $m_{ee}$. It can be seen that the Origin-X is the most promising beyond ton-scale experiment to probe the parameter space of $m_{ee}$ beyond 10 meV for normal ordering.
@@ 5 Leptogenesis within type-I seesaw framework
Leptogenesis has been an elegant framework for generating the observed baryon asymmetry of the Universe, which cannot be addressed within SM. The key idea of leptogenesis is to first create an asymmetry in the lepton sector, and subsequently the created lepton asymmetry becomes a net baryon asymmetry through anomalous $B+L$ violating sphaleron processes [88, 27, 93, 144]. In the original proposal of leptogenesis [74], the CP-violating and lepton number violating out of equilibrium decays of heavy Majorana right-handed neutrinos introduced within the Type-I seesaw mechanism. Many variants of leptogenesis [95, 105, 118, 13, 42] have been explored, and thermal leptogenesis is one of its kind, where the typical mass of right-handed neutrinos is considered at a high scale of around $10^{10}$  GeV or more. Depending on how we account for flavor parts in the determination of CP-asymmetry, there are two classes of thermal leptogenesis: unflavored and flavored type, and we limit ourselves to unflavored leptogenesis in the rest of the analysis.
@@ 6 Conclusion
We have studied a non-supersymmetric version of the Type I seesaw model in which the flavour structure is dictated by non-holomorphic modular forms of level three (polyharmonic Maass forms transforming under $\Gamma_{3}\simeq A_{4}$). The role of modular symmetry is to reduce the model parameters in the Yukawa sector by replacing conventional flavor fields with modular Yukawas. The structure of Yukawa couplings depends only on the complex modulus $\tau$; thereby, the mass matrices are controlled by $\tau$ and a few other minimal parameters. The framework simultaneously explains the observed pattern of lepton mixing and provides calculable predictions for observables that connect low-energy neutrino physics to high-scale leptogenesis.
@@
An important phenomenological implication of the present framework concerns neutrinoless double beta decay ($0\nu\beta\beta$), which serves as a key probe of the Majorana nature of neutrinos and lepton number violation. In our analysis, the estimated effective Majorana mass parameter $m_{ee}$ lies in the sub-meV to few-meV range, $m_{ee}\in[1.9\times 10^{-4},\,5.2\times 10^{-3}]$ eV, for normal ordering. This model predictions on effective Majorana mass parameter and half-lifes are well below the present experimental sensitivities from GERDA [18], KamLAND-Zen [6], and EXO-200 [25] which constrain $m_{ee}\lesssim(0.1-0.4)$ eV as well as the forthcoming ton-scale experiments such as CUPID [26], LEGEND-1000 [9], and nEXO [15], whose projected sensitivities extend down to $\mathcal{O}(0.006-0.027)$ eV. This covers the entire parameter space of $m_{1}-m_{ee}$ for Inverted Ordering and a very thin space in the upper region of the Normal Ordering . However, the limit of $m_{ee}$ that we obtained from our model can be accessed in the future beyond the ton-scale experiments like ORIGIN-X [80], THEIA[84], LEGEND-6000[1] etc as shown in Fig. 9. Moreover, the tight parameter correlations imposed by the modular symmetry ensure that any future measurement of $0\nu\beta\beta$ would place stringent tests on the viability of this model, especially for the normal ordering scenario where the allowed region is already highly constrained.
@@ Acknowledgments
Appendix A Appendix
A.1 Modular $A_{4}$ invariant Lagrangian
@@
Figure 13: We have schematically shown the Fundamental Domain for any data points of $Re[\tau]\;\text{versus}\;Im[\tau]$ satisfying the constraints of a Fundamental Domain in the upper half plane $\mathcal{H}$ .
A.2 Yukawa couplings with different modular weights
@@
$\tau=x+iy,\quad x\equiv\operatorname{Re}[\tau],\quad y\equiv\operatorname{Im}[\tau],\quad q=e^{2\pi i\tau}.$ (A.2)
@@
$\hskip 56.9055pt+\frac{9q^{1/3}}{2\pi}\left(1+\frac{7q}{4}+\frac{8q^{2}}{7}+\frac{9q^{3}}{5}+\frac{14q^{4}}{13}+\frac{31q^{5}}{16}+\frac{20q^{6}}{19}+\dots\right),$ (A.3)
@@
$\hskip 56.9055pt-\frac{81q^{1/3}}{16\pi^{3}}\left(1+\frac{73q}{64}+\frac{344q^{2}}{343}+\frac{567q^{3}}{500}+\frac{20198q^{4}}{2197}+\frac{4681q^{5}}{4096}+\dots\right),$ (A.4)
@@
$M_{\ell}=v\,Y^{(0)}_{\ell}(\tau;\,\alpha_{\ell},\beta_{\ell},\gamma_{\ell}),\qquad M_{D}=v\,Y^{(-2)}_{D}(\tau;\,\alpha_{D},\beta_{D},\gamma_{D}),$ (A.5)
@@
Table 10: Modular Yukawa triplets with three components evaluated at various representative complex moduli parameter $(\tau)$. These Yukawa couplings can be used to construct the charged-lepton and neutrino mass matrices by multiplying each triplet component by the corresponding complex prefactor (e.g. $(\alpha_{\ell},\beta_{\ell},\gamma_{\ell},\alpha_{D},\beta_{D},\gamma_{D})$) as described in the main text. The detailed discussion expansions listed in Appendix A.2.